OpenStudy (anonymous):

Find the slope intercept form for the line satisfying the conditions Perpendicular to y=-1/4x+6 passing through the point (3, -6) I have to find the equation of the line.

OpenStudy (anonymous):

if lines are perpendicular then the product of their slopes is -1 so so -1/4 *4 =-1 so the slope of your line is 4. Next use y-y1=m(x-x1) with the points given eg y- (-6)=4(x-3) and write the answer in y=bx+a form

OpenStudy (akashdeepdeb):

@emma63 Is your query cleared?